Erectors Co. owns earth moving equipment that cost After 8 years it will have estimated salvage value of P18,000.00 Compute the depreciation charge for straight line method for first two years and the book value at the end of 5 years.

1
Expert's answer
2022-02-07T15:34:13-0500

First two years

"Straight\ line\ Depreciation=\frac{(Cost\ of\ Asset - Residual\ value)}{Useful\ life\ of\ an\ Asset}"

"Cost\ of\ Asset=Y"

"Residual\ value=Salvage\ value=18,000"

"Useful\ life=8"

Assuming "Y=100,000"

"Straight line\ Annual\ Depreciation =" "\frac{(100,000-18,000)}{8}" "=10,250"

The equipment depreciates 10,250 annually.

For the first two years,

"\bold{Depreciation = 10,250 \times 2=20,500}"


At end of 5 Years

"Book\ value = Total Cost - Accumulated\ Depreciation"

"Accumulated\ Depreciation=10,250 \times 5=51,250"

"Total\ cost = 100,000"

"\bold{Book\ Value=100,000 -51,250=48,750}"